Assume all future cash flows (FCFF, FCFE, and Dividends) will grow at a constant sustainable growth rate (g) in perpetuity; policy. The companys cost of debt is 20% while the companys cost of equity is 30%; the company has an effective tax rate of 35%. Use the following information in the table below to help answer problems 26-27: nd ide FCFE Today (T-0) FCFF Today (T O) Shareholder Equity Total Debt Total Assets Net Income Dividends Shares Outstanding in millions 800 750 400 600 1000 100 40 25 26. An activist investor who wants to purchase all the companys shares would be willing to pay approximately Do not use the DDM. Use either FCFF or FCFE, whichever is appropriate. a. $165 per share b. $176 per share c. $213 per share d. $230 per share e. $245 per share 27. A private equity group who wants to purchase all of the companys assets would be willing to pay approximately Do not use the DDM. Use either FCFF or FCFE, whichever is appropriate. a. $9,583M b. $10,222M c. $10,518M d. $17,969M e. $19,167M
Answer #1

Answer 26:

Correct answer is:

e. $245 per share

Explanation:

Sustainable Growth rate = Return of Equity * (1 - Dividend payout ratio)

= (Net Income / Shareholder Equity) * (1 - Dividend / Net Income)

= (100/ 400) * (1 - 40 / 100)

= 15%

Sustainable Growth rate = 15%

FCFE today = $800 million

FCFE in year 1 = FCFE today * (1 + Growth rate) = 800 * (1 + 15%) = $920 million

Equity value today (perpetual constant growth) = FCFE in year 1 / (Cost of equity - Growth rate)

= 920 / (30% - 15%)

= $6133.33 million

Share price buyer will be willing to pay = Equity value / Number outstanding shares = 6133.33 / 25 = $245 per share

As such option e is correct and other options a, b, c and d are incorrect

Answer 27:

Correct answer is:

d. $17,969 million

Explanation:

WACC = Cost of equity * Equity proportion + Pretax cost of debt * (1 - Tax rate) * Debt proportion

= 30% * 400 / (400 + 600) + 20% * (1- 35%) * 600 / (400 + 600)

= 19.8%

FCFF today = $750 million

FCFF in year 1 = FCFF today * (1 + Growth rate) = 750 * (1 + 15%) = $862.5 million

Enterprise value today (perpetual constant growth) = FCFF in year 1 / (WACC - Growth rate)

= 862.5 / (19.8% - 15%)

= $17969 million

As such option d is correct and other options a, b, c and e are incorrect